# All-in-one Graph-based Indexing for Hybrid Search on GPUs

### v1

@@ 2.3. Design Principles of Hybrid Index
The preceding analysis reveals that existing retrieval methods face a fundamental trilemma, being unable to simultaneously achieve high accuracy, efficiency, and flexibility. As illustrated in Figure 5, single-path search methods (e.g., HNSW (Malkov and Yashunin, 2018), BM25 (Robertson, 2025), and SPLADE (Formal et al., 2021)) achieve high efficiency but suffer from the semantic gap between vector similarity and end-to-end relevance, resulting in limited accuracy. Multi-path retrieval methods (e.g., Infinity (inf, 2025) and Milvus (mil, 2025a)) enable flexible hybrid search through separate-then-fuse strategies but incur efficiency and accuracy costs, as illustrated previously. Fusion retrieval methods (e.g., IVF-Fusion (Bruch et al., 2024a) and DS-ANN (Zhang et al., 2024)) improve accuracy over single-path approaches while maintaining intermediate efficiency. However, existing approaches are restricted to two-path combinations of dense and sparse vector search, limiting the potential accuracy gains. Moreover, as the combination of dense and sparse vectors is not always effective, they also face concerns about the adaptability to diverse scenarios.
@@
  • Efficiency: Maintaining low-latency retrieval despite multiple paths and heterogeneous distance computations.
Guided by these principles, we propose Allan-Poe, which provides flexible integration of four retrieval paradigms via a well-designed isolated heterogeneous edges mechanism. Unlike separate-then-fuse approaches, Allan-Poe performs path fusion during query processing within a unified index, thereby avoiding the associated accuracy limitations. Additionally, we leverage massive GPU parallelism to achieve both efficient index construction and real-time query performance, meeting the requirement of efficiency.
@@ A USMS is a tuple $H=\{D,F,M_{w}\}$ defined as follows:
  • $M_{w}$: A composite similarity metric $M_{w}:D\times D\rightarrow\mathbb{R}$, defined for any weight vector $w=[w_{d},w_{s},w_{f},w_{k}]\in\mathbb{R}^{4}$ as $M_{w}(q,d)=w_{d}\cdot\text{sim}_{d}(q,d)+w_{s}\cdot\text{sim}_{s}(q,d)+w_{f}\cdot\text{sim}_{f}(q,d)+w_{k}\cdot\text{sim}_{k}(q,d)$, where $\text{sim}_{d}$, $\text{sim}_{s}$, and $\text{sim}_{f}$ denote the inner product similarities, and $\text{sim}_{k}$ represents the path length between query and document entities in the knowledge graph.
@@ 3.2. Hybrid Vector Representation
To integrate dense vector, sparse vector, and full-text retrieval, we employ a vector fusion technique (Wang et al., 2021b) that maps these representations into the USMS. Specifically, dense and sparse vectors are naturally represented in vector form, while keywords used in full-text search can be encoded as sparse vectors where each dimension represents the importance of a term in the vocabulary. For clarity, the sparse vectors generated by full-text search models such as BM25 (Robertson, 2025) are denoted as statistical sparse vectors, while those produced by learning models such as SPLADE (Formal et al., 2021) are called learned sparse vectors. The vector fusion process concatenates these three vector types into a unified high-dimensional representation. Formally, for a document $d$, the concatenated vector is defined as $f_{\text{concat}}(d)=[f_{\text{dense}}(d),f_{\text{sparse}}(d),f_{\text{full}}(d)]$.
@@
Given the RNG-based index constructed from the fused vectors $f_{\text{concat}}(d)$ in UMUS, for any weight vector $w=[w_{d},w_{s},\\ w_{f}]\in\mathbb{R}^{3}$ applied to query vectors $\{f_{\text{dense}}(q),f_{\text{sparse}}(q),f_{\text{full}}(q)\}$, the nearest neighbors can always be retrieved from the index.
@@
We construct a weighted query vector by concatenating the component vectors with their respective weights: $f_{\text{concat}}(q)=[w_{d}*f_{\text{dense}}(d),w_{s}*f_{\text{sparse}}(d),w_{f}*f_{\text{full}}(d)]$. With the Relative Neighborhood Graph (RNG) (Toussaint, 1980) as the index, the nearest neighbors of the vector $f_{\text{concat}}(q)$ can always be found using the greedy search algorithm (Fu et al., 2019; Wang et al., 2021a; Yin et al., 2025). Let $f_{\text{concat}}(d^{*})$ be one of the nearest neighbors. The inner product between the query vector $f_{\text{concat}}(q)$ and candidate vector $f_{\text{concat}}(d^{*})$ expands as $w_{d}*f_{\text{dense}}(q),w_{s}*f_{\text{sparse}}(q),w_{f}*f_{\text{full}}(q)]\cdot[f_{\text{dense}}(d^{*}),f_{\text{sparse}}(d^{*}),f_{\text{full}}(d^{*})]=w_{d}*f_{\text{dense}}(q)\cdot f_{\text{dense}}(d^{*})+w_{d}*f_{\text{sparse}}(q)\cdot f_{\text{sparse}}(d^{*})+w_{f}*f_{\text{full}}(q)\cdot f_{\text{full}}(d^{*})$, which corresponds exactly to the weighted combination of similarities from the three retrieval paths. Consequently, for any weight vector $w=[w_{d},w_{s},w_{f}]\in\mathbb{R}^{3}$, the nearest neighbors of the three types of vectors can always be retrieved. ∎
Theorem 2 establishes that the index supports arbitrary weight combinations without requiring reconstruction. As the overhead of constructing an exact RNG is significant, we construct an approximate RNG utilizing the strategy in NGT (Iwasaki and Miyazaki, 2018) and CAGRA (Ootomo et al., 2024), which has been proven effective and efficient for approximate neighbor retrieval on GPUs (Ootomo et al., 2024).
@@ 3.4. Logical Edges Augmentation
Knowledge graphs can be constructed from the corpus using deep language models such as BERT (Devlin et al., 2019) or LLMs (Fu et al., 2025). For each node in our hybrid index, we store associated entities alongside the fused vector and maintain an entity-to-node mapping. We then extract inter-entity relations from the knowledge graph and represent them as logical edges. Formally, let $V(X)$ denote the entity set of node $X$, and $G(V,R)$ represent the knowledge graph with entities $V$ and relations $R$. The logical edges for node $X$ comprise triplets $\{(s,r,t)\mid s\in V(X),r\in R,t\in V\setminus V(X)\}$, where $s$ and $t$ are entities connected by relation $r$. Thus, any two entities from different nodes that are related in the knowledge graph establish a logical edge between their corresponding document nodes. During search, we dynamically leverage logical edges through a fine-grained entity-document unified strategy that enhances query capability while preserving efficiency (detailed in Section 4.2). Notably, logical edge augmentation is optional, representing a trade-off between potential accuracy improvements and the substantial computational cost of knowledge graph construction (Yang et al., 2025c).
@@ 4.1. Efficient Index Construction on GPU
Allan-Poe’s hybrid index construction presents additional complexity due to the challenges brought by hybrid distance computation and heterogeneous edge establishment. This substantial overhead motivates our use of GPUs to accelerate the construction process. As illustrated in Figure 7, the Allan-Poe indexing pipeline comprises four key procedures.
@@
Procedure 1: Initial $k$-NN Graph Construction with Hybrid Distance. We employ the NN-Descent algorithm (Dong et al., 2011) to construct an approximate $k$-nearest neighbor ($k$-NN) graph from the fused vectors. NN-Descent operates on the principle that ”a neighbor’s neighbors are likely neighbors”—it iteratively refines the graph by evaluating 2-hop neighbors and updating connections based on distance comparisons. This approach has demonstrated superior efficiency to incremental methods on GPUs (Li et al., 2025b; Wang et al., 2021c). Within this framework, hybrid distance calculation between fused vectors represents the most computationally intensive operation, involving two distinct computation patterns. To accelerate this process on GPUs, we assign an entire warp to collaboratively compute each distance. For dense vector computation, each thread fetches four operands using CUDA vectorized instructions (Gale et al., 2020), maximizing memory bandwidth utilization. Threads then multiply their operands in parallel, storing intermediate results in registers. After processing all dense dimensions, the algorithm computes distances for both learned and statistical sparse vectors using identical processing. Sparse vectors are stored in CSR format (Greathouse and Daga, 2014), with non-zero values and indices maintained in separate arrays. Distance calculation is transformed into a parallel set intersection: each thread fetches an index from the candidate’s sparse vector and searches for it in the explored node’s sparse vector using parallel binary search. This design enables the frequently used explored nodes’ vectors to be cached in shared memory, minimizing access overhead. When matches occur, the corresponding values are multiplied and accumulated with the dense vector results. Finally, per-thread values are aggregated using warp-level reduction operations in CUDA.
@@
Procedure 2: RNG-IP Joint Pruning. For the inner product metric, vectors with large norms (lengths) frequently appear in maximum inner product search results, making connections to such vectors an effective strategy for improving search efficiency (Tan et al., 2019). Recent approaches either apply inner product (IP) pruning directly or augment RNG-pruned edges with high-norm nodes in graph indexes (Chen et al., 2025b, c). However, these methods increase index size and often create uneven node degrees, which hinders aligned parallel processing on GPUs. To overcome these limitations, we combine RNG and IP pruning strategies with GPU-specific optimizations, achieving efficient inner product search while preserving both the original index size and aligned degree structure. Our joint pruning approach operates in two phases. The first phase employs RNG pruning from CAGRA to refine neighbors and reduce search complexity. As shown in Figure 7, we calculate the detourable routes for neighbors in the $k$-NN graph in parallel and sort neighbors by their detourable route counts. The second phase applies the IP pruning strategy [^2] For node $A$ with neighbor set $R$, candidate $G$ is excluded if $\exists X\in R$ such that $IP(G,X)>IP(G,G)$. Similarly, $X$ is filtered if $IP(G,X)>IP(X,X)$ (Tan et al., 2019). to remove neighbors with small vector norms. To efficiently achieve the pruning, We parallelize the inner product calculations between the candidate neighbor and current neighbors (e.g., between candidate $G$ and neighbors $X\in\{B,C,D,E,F\}$ in Figure 7), where each warp is responsible for the distance calculation between a current neighbor and the candidate neighbor. Finally, $d/2$ neighbors and $d/2$ reverse neighbors are concatenated to form each node’s final edge list, where $d$ represents the target index degree.
@@
Procedure 3: Keyword-aware Neighbor Recycling. This procedure recycles pruned edges from Procedure 2 according to the strategy established in Section 3.3. A brute-force implementation would incur substantial computational overhead, as checking the condition $K(A)\cap K(X)\subseteq K(Y)$ for each node $A$ and its pruned neighbors $X$ against all current neighbors $Y$ requires numerous set intersection operations. To optimize this process, we leverage computations already performed in previous procedures. Specifically, we assign a Boolean keyword flag to each neighbor of each node. During the second phase of Procedure 2, while computing statistical sparse vector similarities via set intersection between neighbors of node $A$, we check whether non-intersecting keywords exist in $A$’s keyword set $K(A)$. If so, we set the corresponding neighbor’s keyword flag to 1. For efficient $K(A)$ lookups, we maintain it in a hash map within GPU shared memory. For example, in Figure 7, node $E$ is scheduled for pruning during Procedure 2 due to its numerous detourable routes. During the second phase’s distance calculations between $E$ and current neighbors ($B$, $C$, $D$) in Procedure 2, we examine non-intersecting keywords from $E$ (e.g., $K(E)\setminus K(B)=\{k_{1},k_{4}\}$) and verify their presence in $K(A)$ using the hash map. Here, $k_{4}$ exists in $K(A)$ but not in any current neighbor, violating the subset condition. Consequently, $E$’s keyword flag is set to 1. Finally, we traverse all pruned neighbors and recycle those with activated keyword flags as keyword edges.
@@
Procedure 4: Logical Edge Augmentation. In this final procedure, we establish logical edges by mapping knowledge graph entities to their corresponding document nodes in the graph index and creating connections between nodes whose entities share relationships in the knowledge graph.
@@ 17       $KE\leftarrow KE\cup\{v\}$;
19 $SE\leftarrow$ $d/2$ nodes in SESE and $d/2$ reverse neighbors;
@@
4.2. Flexible Query Processing on GPU
4.2.1. Query on the Base Semantic Edges.
As established in Section 3.2, base semantic edges constructed from fused vectors support arbitrary weight combinations across retrieval paths. Given a weight vector $w=[w_{d},w_{s},w_{f}]$ for dense, learned sparse, and statistical sparse vectors respectively, the query vector is formulated as $f_{\text{concat}}(q,w_{d},w_{s},w_{f})=[w_{d}*f_{\text{dense}}(q),w_{s}*f_{\text{sparse}}(q),w_{f}*f_{\text{full}}(q)]$. Single-path retrieval is achieved by setting the corresponding weight to 1 (or any non-zero value) and others to 0. For instance, the fused query vector $f_{concat}(q,1,0,0)$ retrieves documents using only dense vector similarity. This approach extends naturally to two-path and three-path configurations. To optimize search efficiency, we select entry points from nodes with the smallest vector norms. The computationally intensive hybrid distance calculations during query processing are accelerated using the same GPU-optimized strategy described in Section 4.1. The only difference is that during the set intersection operation, each thread fetches an index from the document’s sparse vector and searches for it in the query’s sparse vector. This design reduces time complexity by searching the typically smaller query vector.
4.2.2. Query with Keyword Augmentation.

### v2

@@ 2.3. Design Principles of Hybrid Index
The preceding analysis reveals that existing retrieval methods face a fundamental trilemma, being unable to simultaneously achieve high accuracy, efficiency, and flexibility. As illustrated in Figure 5, single-path search methods (e.g., HNSW (Malkov and Yashunin, 2018), BM25 (Blanco and Boldi, 2012), and SPLADE (Formal et al., 2021)) achieve high efficiency but suffer from the semantic gap between vector similarity and end-to-end relevance, resulting in limited accuracy. Multi-path retrieval methods (e.g., Infinity (inf, 2025) and Milvus (mil, 2025a)) enable flexible hybrid search through separate-then-fuse strategies but incur storage and efficiency costs, as illustrated previously. Fusion retrieval methods (e.g., IVF-Fusion (Bruch et al., 2024a) and DS-ANN (Zhang et al., 2024)) improve accuracy over single-path approaches while maintaining intermediate efficiency. However, existing approaches are restricted to two-path combinations of dense and sparse vector search, limiting the potential accuracy gains. Moreover, as the combination of dense and sparse vectors is not always effective, they also face concerns about the adaptability to diverse scenarios.
@@
  • Efficiency: Maintaining low-latency retrieval and low storage overhead despite multiple paths and heterogeneous distance computations.
Guided by the three principles, we propose Allan-Poe, which provides flexible integration of three retrieval paths augmented by knowledge graphs via a well-designed isolated heterogeneous edges mechanism. Unlike separate-then-fuse approaches, Allan-Poe performs path fusion during query processing within a unified hybrid index, thereby avoiding the associated accuracy limitations. Additionally, we leverage massive GPU parallelism to achieve both efficient index construction and real-time query performance, meeting the requirement of efficiency.
@@ A USMS is a tuple $H=\{D,F,M_{w}\}$ defined as follows:
  • $M_{w}$: A composite similarity metric $M_{w}:D\times D\rightarrow\mathbb{R}$, defined for any weight vector $w=[w_{d},w_{s},w_{f}]\in\mathbb{R}^{3}$ as $M_{w}(q,d)=w_{d}\cdot\text{sim}_{d}(q,d)+w_{s}\cdot\text{sim}_{s}(q,d)+w_{f}\cdot\text{sim}_{f}(q,d)$, where $\text{sim}_{d}$, $\text{sim}_{s}$, and $\text{sim}_{f}$ denote the inner product similarities. Furthermore, $w_{k}\cdot\text{sim}_{k}(q,d)$ is used to augment $M_{w}$ for complex queries, where $\text{sim}_{k}$ represents the path length between query and document entities in the knowledge graph.
@@ 3.2. Hybrid Vector Representation
To integrate dense vector, sparse vector, and full-text retrieval, we employ a vector fusion technique (Wang et al., 2021b) that maps these representations into the USMS. Specifically, dense and sparse vectors are naturally represented in vector form, while keywords used in full-text search can be encoded as sparse vectors where each dimension represents the importance of a term in the vocabulary. For clarity, the sparse vectors generated by full-text search models such as BM25 (Robertson, 2025) are denoted as lexical sparse vectors, while those produced by learning models such as SPLADE (Formal et al., 2021) are called learned sparse vectors. The vector fusion process concatenates these three vector types into a unified high-dimensional representation. Formally, for a document $d$, the concatenated vector is defined as $f_{\text{concat}}(d)=[f_{\text{dense}}(d),f_{\text{sparse}}(d),f_{\text{full}}(d)]$.
@@
The hybrid retrieval problem with any dynamic weight vector $w=[w_{d},w_{s},w_{f}]\in\mathbb{R}^{3}$ is mathematically equivalent to a Maximum Inner Product Search (MIPS) problem in USMS. Consequently, graph-based indices are applicable to the weighted hybrid search without structural modification.
@@
Let the fused vector of a corpus be $f_{\text{concat}}(d)=[w_{d}*f_{\text{dense}}(d),w_{s}*f_{\text{sparse}}(d),w_{f}*f_{\text{full}}(d)]$. The objective of the hybrid search with weight vector $w=[w_{d},w_{s},w_{f}]$ is to maximize the composite similarity score $M_{w}(q,d)$. We construct a transformed query vector $q^{\prime}$ in the Unified Semantic Metric Space by scaling the individual components with their corresponding weights: $q^{\prime}=[w_{d}*f_{\text{dense}}(q),w_{s}*f_{\text{sparse}}(q),w_{f}*f_{\text{full}}(q)]$. Now, we perform the inner product between the transformed query $q^{\prime}$ and the stored corpus vectors $f_{concat}(d)$ in the Unified Semantic Metric Space: $\langle q^{\prime},f_{\text{concat}}(d)\rangle=\langle[w_{d}*f_{\text{dense}}(q),w_{s}*f_{\text{sparse}}(q),w_{f}*f_{\text{full}}(q)],[f_{\text{dense}}(d),f_{\text{sparse}}(d),f_{\text{full}}(d)]\rangle=w_{d}*f_{\text{dense}}(q)\cdot f_{\text{dense}}(d)+w_{d}*f_{\text{sparse}}(q)\cdot f_{\text{sparse}}(d)+w_{f}*f_{\text{full}}(q)\cdot f_{\text{full}}(d)=M_{w}(q,d)$. This derivation proves that finding $d$ that maximizes the weighted hybrid similarity $M_{w}(q,d)$ is exactly equivalent to finding the corpus that maximizes the inner product $\langle q^{\prime},f_{\text{concat}}(d)\rangle$ in USMS. ∎
To achieve efficient maximum inner product search on GPUs, we employ the framework CAGRA (Ootomo et al., 2024) as a backbone of the hybrid index, which has been proven effective and efficient for approximate neighbor retrieval on GPUs (Ootomo et al., 2024).
@@ 3.4. Logical Edges Augmentation
Knowledge graphs can be constructed from the corpus using deep language models such as BERT (Devlin et al., 2019) or LLMs (Fu et al., 2025). For each node in the hybrid index, we store associated entities alongside the vector and maintain an entity-to-node mapping. We then extract inter-entity relations from the knowledge graph and represent them as logical edges. Formally, let $V(X)$ denote the entity set of node $X$, and $G(V,R)$ represent the knowledge graph with entities $V$ and relations $R$. The logical edges for node $X$ comprise triplets $\{(s,r,t)\mid s\in V(X),r\in R,t\in V\setminus V(X)\}$, where $s$ and $t$ are entities connected by relation $r$. Thus, any two entities from different nodes that are related in the knowledge graph establish a logical edge between their corresponding document nodes. During search, we dynamically leverage logical edges through a fine-grained entity-document unified strategy that enhances query capability while preserving efficiency (detailed in Section 4.2). Notably, logical edge augmentation is optional, representing a trade-off between potential accuracy improvements and the substantial computational cost of knowledge graph construction (Yang et al., 2025c). Therefore, selectively applying logical edges to small-scale, high-value data shards serves as a strategic approach to enhance query performance.
@@ 4.1. Efficient Index Construction on GPU
The hybrid index construction presents additional complexity due to the challenges brought by hybrid distance computation and heterogeneous edge establishment. This substantial overhead motivates our use of GPUs to accelerate the construction. As illustrated in Figure 7, the indexing pipeline comprises four key steps.
Figure 7. Index construction process in Allan-Poe.
@@
Step 1: $k$-NN Graph Construction with Hybrid Distance. We employ the NN-Descent algorithm (Dong et al., 2011) to construct an approximate $k$-nearest neighbor ($k$-NN) graph from the fused vectors. NN-Descent operates on the principle that ”a neighbor’s neighbors are likely neighbors”—it iteratively refines the graph by evaluating 2-hop neighbors and updating connections based on distance comparisons. This approach has demonstrated superior efficiency to incremental methods on GPUs (Li et al., 2025b; Wang et al., 2021c). Within this framework, hybrid distance calculation between fused vectors represents the most computationally intensive operation, involving two distinct computation patterns. To accelerate this process on GPUs, we assign an entire warp to compute each distance collaboratively. For dense vector computation, each thread fetches four operands using CUDA vectorized instructions (Gale et al., 2020), maximizing memory bandwidth utilization. Threads then multiply their operands in parallel, storing intermediate results in registers. After processing all dense dimensions, the algorithm computes distances for both learned and lexical sparse vectors using identical processing. Sparse vectors are stored in CSR format (Greathouse and Daga, 2014), with non-zero values and indices maintained in separate arrays. Distance calculation is transformed into a parallel set intersection: each thread fetches an index from the candidate’s sparse vector and searches for it in the explored node’s sparse vector using parallel binary search. This design enables the frequently used explored nodes’ vectors to be cached in the shared memory, minimizing access overhead. When matches occur, the corresponding values are multiplied and accumulated with the dense vector results. Finally, per-thread values are aggregated using warp-level reduction operations in CUDA.
@@
Step 2: RNG-IP Joint Edge Pruning. For the inner product metric, vectors with large norms (lengths) frequently appear in maximum inner product search results, making connections to such vectors an effective strategy for improving search efficiency (Tan et al., 2019). Recent approaches either apply inner product (IP) pruning directly or augment RNG-pruned edges with high-norm nodes in graph indexes (Chen et al., 2025b, c). However, these methods increase index size and often create uneven node degrees, which hinders aligned parallel processing on GPUs. To overcome these limitations, we combine RNG and IP pruning strategies with GPU-specific optimizations, achieving efficient inner product search while preserving both the original index size and aligned degree structure. Our joint pruning approach operates in two phases. The first phase employs RNG pruning from CAGRA to refine neighbors and reduce search complexity. As shown in Figure 7, we calculate the detourable routes for neighbors in the $k$-NN graph in parallel and sort neighbors by their detourable route counts. The second phase applies the IP pruning strategy [^2] For node $A$ with neighbor set $R$, candidate $G$ is excluded if $\exists X\in R$ such that $IP(G,X)>IP(G,G)$. Similarly, $X$ is filtered if $IP(G,X)>IP(X,X)$ (Tan et al., 2019). to remove neighbors with small vector norms. To efficiently achieve the neighbor pruning, We parallelize the inner product distance calculations between the candidate neighbor and current neighbors (e.g., between candidate $G$ and neighbors $X\in\{B,C,D,E,F\}$ in Figure 7), where each warp is responsible for the distance calculation between a current neighbor and the candidate neighbor. Finally, $d/4$ neighbors and $d/4$ reverse neighbors are concatenated to fill each node’s edge list, where $d$ represents the target index degree. For the remaining $d/2$ neighbors, we append the fused semantic neighbors with single-path neighbors. This composite structure effectively approximates the Pareto frontiers, thereby ensuring retrieval robustness under arbitrary weights.
@@
Step 3: Keyword-aware Neighbor Recycling. This procedure recycles pruned edges from Step 2 according to the strategy established in Section 3.3. A brute-force implementation would incur substantial computational overhead, as checking the condition $K(A)\cap K(X)\subseteq K(Y)$ for each node $A$ and its pruned neighbors $X$ against all current neighbors $Y$ requires numerous set intersection operations. To optimize this process, we fully utilize the results calculated in previous steps. Specifically, we assign a Boolean type keyword flag to each neighbor of each node. During the second phase of Step 2, while computing lexical sparse vector similarities via set intersection between neighbors of node $A$, we check whether non-intersecting keywords exist in $A$’s keyword set $K(A)$. If so, we set the corresponding neighbor’s keyword flag to 1. For efficient $K(A)$ lookups, we maintain it in a hash map within GPU shared memory. For example, in Figure 7, node $E$ is scheduled for pruning during Step 2 due to its numerous detourable routes. During the second phase’s distance calculations between $E$ and current neighbors ($B$, $C$, $D$) in Step 2, we examine non-intersecting keywords from $E$ (e.g., $K(E)\setminus K(B)=\{k_{1},k_{4}\}$) and verify their presence in $K(A)$ using the hash map. Here, $k_{4}$ exists in $K(A)$ but not in any current neighbors, violating the subset condition. Consequently, $E$’s keyword flag is set to 1. Finally, we traverse all pruned neighbors and recycle those with activated keyword flags as keyword edges.
@@
Step 4: Logical Edge Augmentation. Finally, we establish logical edges by mapping knowledge graph entities to their corresponding document nodes in the index and creating connections between nodes whose entities share relationships in the knowledge graph.
@@ 17       $KE\leftarrow KE\cup\{v\}$;
19 $SE\leftarrow$ $d/4$ nodes in SESE and $d/4$ reverse neighbors;
20   Append SESE with $d/2$ single path neighbors;